OpenStudy (anonymous):

A worker is filling out his timesheet. He worked 8 hours on Monday, 7 hours and 30 minutes on Tuesday, 8 3/4 hours on Wednesday, 4 hours on Thursday, and 8 1/4 hours on Friday. If he earns $14.35 per hour, what will be his gross pay for this week? I keep getting 520.905

OpenStudy (anonymous):

8+7.30+8.75+4+8.25= X 14.35 = 520.905

OpenStudy (anonymous):

if rounding it off 521.00

OpenStudy (anonymous):

\[\left(8+\left(7+\frac{30}{60}\right)+\left(8+\frac{3}{4}\right)+4+\left(8+\frac{1}{4}\right)\right)*14.35=523.775 \]

OpenStudy (anonymous):

ok you lost me..

OpenStudy (anonymous):

with the fractions..

OpenStudy (anonymous):

Lol. I had some explanation. Then I deleted it... Sigh.

OpenStudy (anonymous):

Let me redo it.

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

\[\left(8+\left(7+\frac{30}{60}\right)+\left(8+\frac{3}{4}\right)+4+\left(8+\frac{1}{4}\right)\right)=\left(8+\frac{15}{2}+\frac{35}{4}+4+\frac{33}{4}\right)=\frac{73}{2}=36.5 \]\[36.5*14.35=523.775 \]

OpenStudy (anonymous):

First, find the total hours he works for the week. \[8+7{1\over2}+8{3\over4}+4+8{1\over4}\] \[35+{1\over2}+{3\over4}+{1\over4}\] \[35+{2\over4}+{3\over4}+{1\over4}\] \[35+{6\over4}\] \[35+1{2\over4}\] \[36+{1\over2}\] \[14.35(36+{1\over2})\]

OpenStudy (anonymous):

Can you solve the final one?

OpenStudy (anonymous):

okkay!i think i got it

OpenStudy (anonymous):

holdon

OpenStudy (anonymous):

Basically it's just following the formula 14.35h (pay x hours)

OpenStudy (anonymous):

yeaa.. i got it now

OpenStudy (anonymous):

thanks again for your help. =)

OpenStudy (anonymous):

:) good.

OpenStudy (anonymous):

Just remember that 7.30 is \[7{1 \over 2} hours\] It's much easier to make time into a fraction and add it all up rather than converting time into standard unit...

OpenStudy (anonymous):

okay, thanks =)

OpenStudy (anonymous):

@monokerous Shouldn't it be: \[7:30=7\frac{1}{2}\]

OpenStudy (anonymous):

That's what I did.

OpenStudy (anonymous):

Not a big deal, however, you wrote "Just remember that 7.30 is . . . "

OpenStudy (anonymous):

oh. Right. My bad. 7:30. Minor detail. :/ Thanks for that!

OpenStudy (anonymous):

np